OpenStudy (anonymous):

Find f(5) for this sequence: f(1)=2 and f(2)=3, f(n)= f(n-1)+f(n-2), for n>2

OpenStudy (anonymous):

start with n=3

OpenStudy (anonymous):

I don't know how to do it because I am checking my sons homework

OpenStudy (anonymous):

equation is given put n=3 u will get f(3)...continue..same...u will get f(4) ..then f(5)

OpenStudy (anonymous):

by the logic: f(n) = f(n -1) + f(n- 2) for n > 2 f(3) = f(3-1) + f(3 - 2) which is: f(3) = f(2) + f(1) f(3) = 3 + 2 (given in problem) f(3) = 5 using the sequence again: f(4) = f(4 - 1) + f(4 - 2) f(4) = f(3) + f(2) f(4) = 5 + 3 f(4) = 8 do you want to try and find f(5)?

OpenStudy (anonymous):

yes u r going right ...now calculate f(5)

OpenStudy (anonymous):

f(5) = f(4) + f(3)